OpenStudy (anonymous):

What is the slope of the line that is parallel to the line whose equation is 2x + y = 4. Type a numerical answer in the space provided. Do not include spaces in your answer. If necessary, use the / key to represent a fraction bar and only type improper fractions (no mixed numbers).

OpenStudy (texaschic101):

a parallel line will contain the same slope as the slope in the given equation. 2x + y = 4 To find the slope, put your equation in y = mx + b form, where m is your slope. So put your equation in y = mx + b form...give it a try

OpenStudy (texaschic101):

hint : subtract 2x from both sides

OpenStudy (anonymous):

ok is that it

OpenStudy (anonymous):

so y=4-2x

OpenStudy (texaschic101):

no.. 2x + y = 4 -- subtract 2x from both sides y = -2x + 4 as you can see, the number in the m position is -2. That is your slope

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

your slope is -2

OpenStudy (anonymous):

ok is that the answer to the problem

OpenStudy (phi):

fyi @thatpersonthatneeds so y=4-2x is correct, but by convention people write it in the other order y = -2x + 4 this is so it is in the form y = mx + b the m (number in front of the x) is the slope. (that is why people write the equation this way... you can read off the slope by looking for the number in front of the x) the y intercept is the b in your problem y = -2x + 4 -2 is the slope and 4 is the y-intercept

OpenStudy (anonymous):

ok